OpenStudy (anonymous):

350 is 70% of what number? A) 500 B) 420 C) 245 D) 280 What percent of 64 is 24? A) 267% B) 62.5% C) 40% D) 37.5% What percent of $6.50 is $2.60? A) 25% B) 40% C) 90% D) 250% 20% of 140 is what number? A) 280 B) 120 C) 28 D) 12

OpenStudy (anonymous):

a) 350 / 70% b) 64 * x%=24 x%=24/64 c) same approach with b d) ummmm....i am also confused by the wordings..

OpenStudy (anonymous):

d) 140*20%??? @texaschic101

OpenStudy (texaschic101):

All of these problem can be solved using the percentage formula : is/of = %/100 350 is 70% of what number is = 350 % = 70 of = x now sub 350/x = 70/100 --- cross multiply 70x = 35000 x = 35000/70 x = 500 so 350 is 70% of 500 check... 350 = 0.70(500) 350 = 350 (correct) Your other percentage problems can be solved the same way...just be sure you put your numbers in the right places. You wanna try the next one ?

OpenStudy (texaschic101):

confused with D......use the formula is/of = %/100 20% of 140 is what number % = 20 of = 140 is = x now sub x/140 = 20/100 -- cross multiply 100x = (20)(140) 100x = 2800 x = 2800/100 x = 28 check.. 20% of 140 is 28 0.20 (140) = 28 28 = 28 (correct) so 20% of 140 = 28